### 15 Dec 2023 v2

@@ 3.3.2 Visualization Analysis
We further performed a visual analysis to measure the performance of Gene-MOE. We randomly selected 1,000 patients from the test set to perform this evaluation. The real gene expression of test patients was first fed into Gene-MOE to generate the reconstruction expression. Then, we used TSNE and UMAP to perform the visualization and evaluated the similarity of these two distributions. Figure S3(c) shows the visualization result using these two methods. According to Figure S3(c), the real gene distribution perfectly coincided with the reconstructed gene distribution, which proves that the Gene-MOE model can perform reconstruction of the input genes more accurately based on the input real genes. Furthermore, Gene-MOE completed the reconstruction based on the low-dimensional feature obtained by the backbone model, which further reinforces that Gene-MOE can learn rich feature representation by the backbone network.
@@
Table 2: Comparison of Concordance Index on 14 cancer types using four distinct models
[…]
@@
[…]
Table 3: Comparison of classification metrics using four distinct models
[…]
@@
[…]
3.4 Ablation Studies
In this section, we extend our analysis by presenting ablation experiments conducted to evaluate the performance of the model proposed in this paper. Four distinct models were constructed for this purpose: 1) the baseline model comprising two FCN layers, 2) the model incorporating MOE by replacing the FCN layers, 3) the Gene-MOE model, and 4) the pre-trained Gene-MOE model. Subsequently, survival analysis and cancer classification tasks were performed using these four models. The results of these tasks are presented in Tables 2 and 3.
@@ 3.4 Ablation Studies
3.4.1 Performance of MOE
By comparing survival analysis result in Table 2, we found the model with the MOE layer could achieve better Concordance results than baseline on 12 cancer types. Moreover, in Table 3, the model with MOE showed better accuracy, and F1-score. These results indicate that using the MOE layer can increase the performance of genomic analysis, which proves our theory that using MOE can make a model learn rich features during the feature extraction process.
3.4.2 Performance of MOAE
By comparing the Concordance Index for model with MOE and the Gene-MOE model in Table 2, we observed that Gene-MOE outperformed the model with MOE on nine cancer types. Furthermore, in Table 2, we found that Gene-MOE performed better in accuracy, recall, precision, and F1-score. These findings prove that the MOAE model can further improve the accuracy of the model by improving the ability to learn deep semantic correlated features.
3.4.3 Performance of Pre-training
By comparing the performance of the Gene-MOE and pre-trained Gene-MOE models in Table 2, we found that the pre-trained Gene-MOE model performed better in eight cancer types. At the same time, the Gene-MOE model demonstrated the same performance on KIRC and LGG datasets. These results reveal the effectiveness of pre-training, particularly in enhancing the performance of most survival analysis models. Further comparison of Gene-MOE and pre-trained Gene-MOE in Table 3 showed that that the pre-trained Gene-MOE model performed better in the precision and recall, but the accuracy and F1-score were lower than that of Gene-MOE. By further compare the F1-score of these two model in Figure 6, we found that the pre-trained model performed less effective on cancer types with few samples, such as the READ dataset. This discrepancy can be attributed to the fact that the pre-training stage primarily captures common cancer features, potentially missing certain characteristics for cancer types with small sample sizes. Consequently, during the fine-tuning stage, the model tends to favor learning patterns from cancer types with larger sample sizes.
4 Conclusion
@@
5 Data Availability
Data in our experiments are available at https://xenabrowser.n
et/datapages/. Codes in our experiments are available at https://github.com/menggerSherry/gene-moe.
6 Funding

### 18 Dec 2023 v3

@@ III-C2 Visualization Analysis
We further performed a visual analysis to measure the performance of Gene-MOE. We randomly selected 1,000 patients from the test set to perform this evaluation. The real gene expression of test patients was first fed into Gene-MOE to generate the reconstruction expression. Then, we used TSNE and UMAP to perform the visualization and evaluated the similarity of these two distributions. Figure 8(c) shows the visualization result using these two methods. According to Figure 8(c), the real gene distribution perfectly coincided with the reconstructed gene distribution, which proves that the Gene-MOE model can perform reconstruction of the input genes more accurately based on the input real genes. Furthermore, Gene-MOE completed the reconstruction based on the low-dimensional feature obtained by the backbone model, which further reinforces that Gene-MOE can learn rich feature representation by the backbone network.
@@
TABLE II: Comparison of Concordance Index on 14 cancer types using four distinct models
BLCA BRCA HNSC KIRC LGG LIHC LUAD LUSC OV SARC CESC STAD COAD UCEC
Baseline 0.643 0.692 0.632 0.728 0.857 0.691 0.649 0.558 0.590 0.722 0.720 0.607 0.674 0.699
@@
TABLE III: Comparison of classification metrics using four distinct models
Accuracy Precision Recall F1-Score
Baseline 0.9304 0.9235 0.8721 0.8799
@@
III-D Ablation Studies
In this section, we extend our analysis by presenting ablation experiments conducted to evaluate the performance of the model proposed in this paper. Four distinct models were constructed for this purpose: 1) the baseline model comprising two FCN layers, 2) the model incorporating MOE by replacing the FCN layers, 3) the Gene-MOE model, and 4) the pre-trained Gene-MOE model. Subsequently, survival analysis and cancer classification tasks were performed using these four models. The results of these tasks are presented in Tables II and III.
@@ III-D Ablation Studies
III-D1 Performance of MOE
By comparing survival analysis result in Table II, we found the model with the MOE layer could achieve better Concordance results than baseline on 12 cancer types. Moreover, in Table III, the model with MOE showed better accuracy, and F1-score. These results indicate that using the MOE layer can increase the performance of genomic analysis, which proves our theory that using MOE can make a model learn rich features during the feature extraction process.
III-D2 Performance of MOAE
By comparing the Concordance Index for model with MOE and the Gene-MOE model in Table II, we observed that Gene-MOE outperformed the model with MOE on nine cancer types. Furthermore, in Table II, we found that Gene-MOE performed better in accuracy, recall, precision, and F1-score. These findings prove that the MOAE model can further improve the accuracy of the model by improving the ability to learn deep semantic correlated features.
III-D3 Performance of Pre-training
By comparing the performance of the Gene-MOE and pre-trained Gene-MOE models in Table II, we found that the pre-trained Gene-MOE model performed better in eight cancer types. At the same time, the Gene-MOE model demonstrated the same performance on KIRC and LGG datasets. These results reveal the effectiveness of pre-training, particularly in enhancing the performance of most survival analysis models. Further comparison of Gene-MOE and pre-trained Gene-MOE in Table III showed that that the pre-trained Gene-MOE model performed better in the precision and recall, but the accuracy and F1-score were lower than that of Gene-MOE. By further compare the F1-score of these two model in Figure 9, we found that the pre-trained model performed less effective on cancer types with few samples, such as the READ dataset. This discrepancy can be attributed to the fact that the pre-training stage primarily captures common cancer features, potentially missing certain characteristics for cancer types with small sample sizes. Consequently, during the fine-tuning stage, the model tends to favor learning patterns from cancer types with larger sample sizes.
IV Conclusion
@@
V Acknowledgment